MICROWAVE AND AB INIT IO STUDIES OF THE XE-CH_{4} VAN DER WAALS COMPLEX

\underline{QING WEN} AND WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB T6G 2G2, Canada.

We previously reported on experimental and theoretical studies of the Xe-N_{2} complex as a test case to investigate Xemolecule interaction potentials.^{a} Here, we present our work on the Xe-CH_{4} complex, which is another appealing test case because of the high symmetry of methane monomer and its anticipated near free internal rotation. As a starting point, an ab initio potential energy surface was constructed at the CCSD(T) level of theory. The recently developed aug-cc-pVQZ-PP basis set was used for the xenon atom and the aug-cc-pVTZ basis set for the other atoms. The basis sets were supplemented with bond functions. Dipole moments were also calculated at various configurations. The reliability of the ab initio calculations was tested against the experimental rotational transition frequencies. Rotational transitions were measured using a pulsed-nozzle Fourier transform microwave spectrometer. Transitions within three internal rotation states, namely the j=0, K=0; j=1, K=0; and j=2, K=1 states, were measured and assigned. Nuclear quadrupole hyperfine structures due to the presence of ^{131}Xe (I=3/2) were detected and analyzed. It was found that the j=1, K=0 state is perturbed by a Coriolis interaction with a nearby j=1, K=1 state. For isotopomers containing CH_{3}D and CHD_{3}, the j=2 states are no longer metastable and could not be observed. The spectroscopic results were used to derive information about the molecular structure and intermolecular dynamics of the Xe-CH_{4} complex.

aQ. Wen and W. Jäger, paper WH09, 59^{th} International Symposium on Molecular Spectroscopy, Ohio State University, Columbus, Ohio, USA, June 21 − 25, 2004.